APACHE POINT OBSERVATORY SDSS 2.5M OBSERVING LOG Thursday December 21, 2006 (MJD 54091) ---=== OBSERVING TEAM ===--- Howard Brewington (Days) Dan Long and Kaike Pan (Observers) Visitor: Shi HuoMing ---=== OBSERVING PLAN ===--- Science! ---=== OBSERVING SUMMARY ===--- Due to blowing snow, we never open tonight. We did part of MSP imager calibraion summarized as: QeCCD r 1 is Run 6635 frames 2 - 49 QeCCD r 2 is Run 6635 frames 50 - 96 QeCCD r 3 is Run 6638 frames 2 - 36* NOTE: endNight complained about untransferring gangs for run 6640. It was a useless run (Dan had trouble when taking post MSP calibration bias, see log). ---=== OBSERVING LOG ===--- Afternoon: ------------- The PT became unhappy during the dome flats; see PT log for more details. The spectro focus was checked and adjusted (by -5000) to the temp of about -4 C. See Problems Section for the focus numbers. I had a problem when opening a tccMon session and could not get show scale or show focus updates for example. Night: ------------- 01:00Z We had blowing snow, so we could not open. We closed doors, and setup for MSP imager calibration. When took measurements, watcher complained about archiving (this is due to the known checksum problem) for MSP calibration frames. Calibrator Summary: QeCCD r 1 is Run 6635 frames 2 - 49 QeCCD r 2 is Run 6635 frames 50 - 96 QeCCD r 3 is Run 6638 frames 2 - 36* *incomplete Ended with lambda 680, time 10 Calibrator Run Notes: ---------------------- Warning during r1 sequence: --------------------------- getFrameMSP 6635 36 11 /home/observer/calibdata/idR-6635-r1-36.fit completed --------------------------------------------------------------------- Warning! too low flux on photodiode! aflux 0. < 5000*(exptime) getFrameMSP 6635 37 11 /home/observer/calibdata/idR-6635-r1-37.fit completed --------------------------------------------------------------------- WARNING! We receive flux too low even after a retry! MSP: Unsolicited response "flux 271.4 aflux 2646080.2 volts 0.0000 avolts -0.2190 voltsRms 0.0008 avoltsRms 0.0013 " ping error 4 stageX 23100 lambda 686.0 time 50.000 6635 38 r1 im cameraNag {} t-g-sdss-2.apo.nmsu.edu 3500 2006-12-22 03:12:46Z: opening camera doors AD1 error from VME as well "Dead" there was definitely light on the chip and the flux level from the "unsolicited response" looks reasonable. Maybe since it got the response at a time when it wasn't expecting it it considered aflux to be zero? QE r2 completed without error. QE r3 ran until frame 37 iop hung during frame 37 exposure at the line "There is already a teamster..." could not get prompt back. had to kill process from another window to get imager control back. Subsequent attempts at either goDrift or goStare would hang im> goDrift -bias camera not setup for drifting, performing an initData statusEntry daqDrifting on node id1 is not 1 but 0 DA not setup for drifting, performing an initData I could still talk to the imager and get imstat returns etc. No vme or DA errors being reported. daqBusy was low. A fourth try at goDrift worked. What changed? Got started again with staring bias. set iccd 3 set color r set msprow [getrow $color] getIllumCol 3 temp_r3.list (edited to take up where previous run crashed) It started taking data with stageX set to 3, when it needed to be 205100. Tried setting msprow to 205100, still got stageX of 3. Snooped around code to find where stageX was set. It appeared to use a variable $column, so I set column to 205100. Still got stageX of 3. So much for starting in the middle. There isn't enough time left in the night to start an entire chip sequence, so shutting things down here at 10:50Z. 06:20Z Although much less than hours ago, we still saw blowing snow. But we feel that PT dome position is high enough to open safely. We opened it and try to get some manual target and secondaries (sky is photometric). 10:55Z post MSP calibration bias. 12:05Z endNight finished. endNight did complained about untransfering gangs for run 6640. 6640 is a run when we got trouble for imager bias (see log), it was a useless run. ---=== IMAGING RUN SUMMARY ===--- Run Time Stripe Lambda Last Flavor Comments Start End Begin End Frame ------------------------------------------------------------------------- 6634 21:31Z 21:50Z 100 O 122.39 127.10 46 ignore 6636 07:09Z 07:12Z 100 O -92.72 -92.13 17 ignore 6639 10:07Z 10:10Z 100 O -48.25 -47.49 19 ignore 6641 10:52Z 10:57Z 100 O -37.04 -35.75 23 ignore ---=== IMAGING RUN DETAILS ===--- ---=== SKIPPY RESULTS ===--- Run Frame nFrames stars muErr muRms nuErr nuRms rot az el --------------------------------------------------------------------------- ---=== LTMATCH RESULTS ===--- Run Field nFields alt az nGood rowMean rowSig colMean colSig rot ------------------------------------------------------------------------ ---=== SPECTROSCOPY DATA SUMMARY ===--- Summary Checked (y/n): QA Procedures Done (y/n): UT Exp Time flavor comment (S/N)^2 totals ========================================== b1 r1 b2 r2 ----- sequence 42072, plate -9999 ------- 21:36 42072 0.0 bias ----- sequence 42073, plate 2580 ------- 22:40 42073 10.0 flat 22:42 42074 2.0 arc 22:49 42075 3.0 arc 22:52 42076 2.0 arc ----- sequence 42077, plate -9999 ------- 04:51 42077 0.0 bias ----- sequence 42078, plate -9999 ------- 06:46 42078 0.0 bias ---=== TELESCOPE OFFSETS AND SCALE I ===--- Time Instrument Az Alt Rot Scale pos offset pos offset pos offset ------------------------------------------------------------------------------ ---=== TELESCOPE OFFSETS AND SCALE II ===--- ---=== FOCUS LOG ===--- setmir piston Temp Wind Time Inst scale M1 M2 Foc Az Alt (C) MPH Dir filt fwhm ------------------------------------------------------------------------------ ---=== WEATHER LOG ===--- Wind Time Temp F Dewp F MPH Direction Dust DIMM Sky 16:56Z 19 12 10 344 (NNW) 153 - clear 17:28Z 20 13 10 23 (NNE) 174 - 18:00Z 20 13 9 358 (NNW) 198 - 18:32Z 22 12 7 311 (NW) 165 - 19:04Z 22 14 7 7 (N) 155 - 19:36Z 23 16 2 353 (NNW) 152 - 20:09Z 22 15 6 350 (NNW) 207 - 20:41Z 24 16 10 0 (N) 169 - 21:13Z 25 14 7 10 (N) 246 - 21:43Z 25 17 8 15 (NNE) 202 - 22:14Z 24 18 9 24 (NNE) 218 - 22:44Z 23 17 10 9 (N) 241 - 23:17Z 23 16 16 31 (NNE) 219 - 23:49Z 23 13 10 31 (NNE) 137 - 00:38Z 23 13 14 25 (NNE) 136 - 01:10Z 23 13 14 12 (NNE) 144 - 01:46Z 24 13 16 10 (N) 139 - 02:18Z 24 14 23 19 (NNE) 124 - 02:51Z 25 14 16 14 (NNE) 165 - 03:23Z 25 15 25 23 (NNE) 123 - 03:55Z 25 15 17 19 (NNE) 199 - 04:27Z 27 15 17 31 (NNE) 119 - 04:59Z 29 13 15 15 (NNE) 111 - 05:32Z 28 14 2 354 (NNW) 75 - 06:04Z 27 15 11 36 (NE) 62 - 06:34Z 28 14 9 355 (NNW) 75 - clear 07:06Z 29 12 11 351 (NNW) 31 - 07:38Z 28 11 13 348 (NNW) 40 - 08:08Z 28 14 6 1 (N) 22 - 08:40Z 30 12 4 9 (N) 53 - 09:10Z 31 12 12 342 (NNW) 36 - 09:41Z 29 12 15 336 (NNW) 58 - 10:12Z 30 12 14 347 (NNW) 64 - 10:43Z 29 12 13 346 (NNW) 49 - 11:13Z 28 13 9 11 (N) 50 - 11:45Z 28 13 10 338 (NNW) 48 - ---=== TELESCOPE STATUS ===--- 23:20Z Door and louvers opened; fans on. Due to blowing snow, we never open. Stowed at 30 deg alt, Cartridge 4 installed. Spectro auto-fill connected. Spectro dewar 11:55Z Dewar wt 78 lb Imager dewar 11:55Z Dewar wt 194 lb ---=== SOFTWARE USED ===--- IOP/SOP: v4_29_0 Watcher: v2_35_0 MCP: v6_3_2 TPM: tpm_v3_5_0 AstroDa: v15_13 TCC: TCC 2.7.2.1 August 6 2004 sdssProcedures: SoS: v5_1_4 hoggPT: v1_6_9 plate-mapper: v4_3_1 ---=== MIRROR NUMBERS ===--- PRIMARY: -------- Scale: 1.000000 MIGS TONIGHT NOMINAL Axial A 24.6750 24.6730 Axial B 26.2700 26.2680 Axial C 27.9330 27.9320 Trans D 13.5830 13.5820 Lateral E 9.9695 9.9695 Lateral F 11.5570 11.5443 GALILS Commanded: 4600. -2000. 2650. -9300. -9800. -9800. Actual: 4600. -2000. 2650. -9300. -9800. -9800. SETMIR VALUES PriDesOrient: 0.00 0.00 0.00 0.00 0.00 PriOrient: 0.00 0.00 0.00 0.00 0.00 SECONDARY: ---------- Focus: 0.00 Air Temp.: -3.1 Alt.: 30.274180 MIGS TONIGHT NOMINAL Axial A 1.5550 1.5540 Axial B 1.1380 1.1390 Axial C 1.1910 1.1920 Trans D -0.0480 -0.0510 GALILS Commanded: 1614043. 1593400. 1543101. -10550. -5800. Actual: 1614050. 1594700. 1543100. -10550. -5800. SETMIR VALUES SecDesOrient: 1257.00 20.00 -10.00 -100.00 176.64 SecOrient: 1256.65 19.69 -10.34 -100.66 176.76 ---=== PROBLEMS IN DETAIL ===--- --== Spectro Focus Numbers ==-- IDL 5.5a> collimate, 42075 Reading /data/spectro/54091/sdR-b1-00042075.fit speclog_version: Command not found. Reading /data/spectro/54091/sdR-b1-00042076.fit speclog_version: Command not found. COLLIMATE: Min offset = -1.21 pix COLLIMATE: Max offset = 0.94 pix COLLIMATE: RMS across CCD = 0.52 pix COLLIMATE: Mean offset = -0.11 pix COLLIMATE: Camera b1 appears to be IN-FOCUS (|mean| < 0.15 pix) COLLIMATE: Predict blue camera ring movement of 3.9 degrees for spectro-1 COLLIMATE: (if red is already in focus) Log file = Collimate-54091-b1-00042075.log PostScript file = Collimate-54091-b1-00042075.ps Reading /data/spectro/54091/sdR-b2-00042075.fit speclog_version: Command not found. Reading /data/spectro/54091/sdR-b2-00042076.fit speclog_version: Command not found. COLLIMATE: Min offset = -0.55 pix COLLIMATE: Max offset = 0.06 pix COLLIMATE: RMS across CCD = 0.11 pix COLLIMATE: Mean offset = -0.08 pix COLLIMATE: Camera b2 appears to be IN-FOCUS (|mean| < 0.15 pix) COLLIMATE: Predict blue camera ring movement of 2.8 degrees for spectro-2 COLLIMATE: (if red is already in focus) Log file = Collimate-54091-b2-00042075.log PostScript file = Collimate-54091-b2-00042075.ps Reading /data/spectro/54091/sdR-r1-00042075.fit speclog_version: Command not found. Reading /data/spectro/54091/sdR-r1-00042076.fit speclog_version: Command not found. COLLIMATE: Min offset = -0.65 pix COLLIMATE: Max offset = 0.41 pix COLLIMATE: RMS across CCD = 0.28 pix COLLIMATE: Mean offset = -0.16 pix COLLIMATE: Camera r1 appears to be OUT-OF-FOCUS (|mean| > 0.15 pix) COLLIMATE: Predict (red) piston movement of 1445 steps for spectro-1 COLLIMATE: Issue SOP command: "sp1; mechPiston 1445" Log file = Collimate-54091-r1-00042075.log PostScript file = Collimate-54091-r1-00042075.ps Reading /data/spectro/54091/sdR-r2-00042075.fit speclog_version: Command not found. Reading /data/spectro/54091/sdR-r2-00042076.fit speclog_version: Command not found. COLLIMATE: Min offset = -0.60 pix COLLIMATE: Max offset = 0.08 pix COLLIMATE: RMS across CCD = 0.16 pix COLLIMATE: Mean offset = -0.16 pix COLLIMATE: Camera r2 appears to be OUT-OF-FOCUS (|mean| > 0.15 pix) COLLIMATE: Predict (red) piston movement of 1457 steps for spectro-2 COLLIMATE: Issue SOP command: "sp2; mechPiston 1457" Log file = Collimate-54091-r2-00042075.log PostScript file = Collimate-54091-r2-00042075.ps % Program caused arithmetic error: Floating illegal operand IDL 5.5a>